If 15 men can do a piece of work in 90 days, in how many days can 30 men do the same work?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If 15 men can do a piece of work in 90 days, in how many days can 30 men do the same work?
-------------
Number of men and days to do the job are inversely related.
-------------
Equation:
days = k/number of men
---
Solve for "k":
90 = k/15
k = 15*90 = 1350
-----
Equation for your problem:
days = 1350/number of men
------
Your Problem: in how many days can 30 men do the same work?
days = 1350/30
days = 45
=================
Cheers,
Stan H.

Answer by richwmiller(17219)

We are assuming that the other 15 added will work at the same pace/rate as the first 15 which is not always a wise assumption.
in one day 15 men do 1/90 of the job
so 30 men should be able to do 2/90 in the same day or 1/45 of the job in one day
so in 45 days 30 men should be able to finish the job